Derek the contractor took out a loan to purchase lumber forone of his projects. He ended up 5328.85 in interestover an 11 month period it theaterestrate on the loan was9.5% how much dhe end up spending on lumber?What was the principal amount of the loan?

1
Expert's answer
2021-06-17T11:35:59-0400

Monthly payment×(Number of months"-" Loan) =Interest


=>Monthly payment×n"-" Loan=328.85


=>"Loan×\frac {(\frac {R} {12} )} {(1-{1} {(1+\frac {R} {12} )^n)}} ×n-Loan=328.85"


"=P×\frac {(\frac {9.5} {12} )} {(1-{1} {(1+\frac {9.5} {12} )^{11} )}} ×11-1=328.85"


Principal Amount =6833.36